sadie5:

Which equation is true for the value b = 10? A. 2(b + 4) = 16 B. 2(b + 2) = 40 C. 3(b – 2) = 24 D. 2(8 + b) = 42 E. 3(b – 4) = 20

sadie5:

do you know this

Joe348:

I believe its C

snowflake0531:

b=10 2(b + 4) = 16 2(b + 2) = 40 3(b – 2) = 24 2(8 + b) = 42 3(b – 4) = 20 Do you know the distributive property?

sadie5:

i think its A

snowflake0531:

funny how you pick all the ones that are false xdd

sadie5:

xd

FredrickStar:

i agree i think it's c

snowflake0531:

shoot kek

Joe348:

thank you^